2x + 3y + 3z = 6
-x + y + z = 2

Answer by robertb(5830)

Multiply the bottom equation by 2.
-2x+2y+2z = 4. Add corresponding sides with the top equation, to get
5y +5z = 10, or y+z = 2. Substitute this into the bottom equation:
-x + (2-z) +z = 2, or -x+2 = 2, or -x = 0, or x = 0. However, y is still y = 2-z, and so the system is dependent, or has infinitely-many ordered triples as solution. The solution set is the set of all triples (0,2-z,z), where z is any real number.